1. An airplane is flying at a constant speed. The thrust from the propeller is 8 000 N [E]. A crosswind is applying a force of 2 000 N [N]. What is the net force acting on the airplane?
1
Expert's answer
2021-07-08T09:57:01-0400

The net force is


"F_\text{net}=\sqrt{8000^2+2000^2}=8246\text{ N}."

It is directed

"\theta=\arctan\frac{ 8000}{2000}=76°"

N of E.